Skip to end of metadata
Go to start of metadata

Kurt and Dan connected v04 to power, I visited later to get it on the net and check the data.

Ubiquiti

  • Turned off frequency list.
  • Fixed AP MAC, replaced current setting with tnw10 WSW: 44:D9:E7:C2:D5:A8.
  • Changed name to Vall_04.

Eventually it linked successfully, signal at -75 dbm, and the dsm could ping the Internet.

DSM

  • KZRAD in port 6 looks like a mote binary message.
  • 12.8V on one mote (probably soil?), 13.3V on the other
  • I see mote IDs in the 0x54 to 0x6b range, so I think KZRAD is working.
  • 2m sonic on port 1 looks good.
  • 10m winds on port 5 look good (repo config still had port 2).
  • USB archive at 5% and growing.

At first there was nothing for 10m LICOR on port 3. Port 3 has 3A fuse and the fuse is good. Port 3 is jumpered for 232. When I checked the fuse that cycled the power on the LICOR, and that seems to have fixed it...

Cannot update eol packages, apparently because somehow some of the directories in the EOL debian ftp archive have become inaccessible. So until I can fix that, I updated the DSM manually by installing xinetd and check-mk, and downloaded and installed the latest package files for isfs-field-erpdigao, eol-daq, and isfs-daq.

The repo config has been fixed to match what was running on the DSM.